OpenStudy (anonymous):

please help to solve and i will fan and give a medal Y" + x y' + y= 0 when y(0)=1 and y'(0)=0

OpenStudy (anonymous):

then just plug in the variables, and its just system of equations so when y(0), it becomes \[y"+xy'+y=0\] For y(0), \[y''+xy'+0=1\] which means \[y''+xy'=1\] then for y'(0) \[y''+x(0)+y=0\] Anything multiplied by 0 is 0, so that means \[y''+0+y=0\] put y on the other side, \[y''=-y\] and you can go plug that in for the other equation, y(0)=1 \[(-y)+xy'+(0)=1\] \[(0)+xy'+(0)=1\] \[xy'=1\] Since when you put 0 in for y', it becomes 0, that means x=1

OpenStudy (anonymous):

does that mean this is another process of solving differential equation ivp ?